Your company has earnings per share of $3.96 . It has 1.2 million shares outstanding, each of which has a price of $48. You are thinking of buying TargetCo, which
has earnings per share of $0.99, 1.4 million shares outstanding, and a price per share of $25. You will pay for TargetCo by issuing new shares. There are no
expected synergies from the transaction.
  
a. If you pay no premium to buy TargetCo, what will your earnings per share be after the merger?
The EPS after the merger is $______ . (Round to the nearest cent.)


b. Suppose you offer an exchange ratio such that, at current pre-announcement share prices for both firms, the offer represents a $20% premium to buy TargetCo.
What will your earnings per share be after the merger?

The EPS after the merger is $_____ . (Round to the nearest cent.)


c. What explains the change in earnings per share in part (a )? (Select the best choice below.)
A. EPS always decline if the firm issues new shares to pay for a merger.
B. EPS declines because you are over − paying for TargetCo.
C. EPS declines because TargetCo has a higher price-earnings ratio than your firm.


Are your shareholders any better or worse off? (Select the best choice below.)
A. In this case, your shareholders are worse off.
B. In this case, your shareholders are better off.
C. In this case, your shareholders are neither worse nor better off.


d. What will your price-earnings ratio be after the merger (if you pay no premium)? How does this compare to your P/E ratio before the merger? How does this
compare to TargetCo's premerger P/E ratio?
The P/E ratio after the merger is_______ . (Round to two decimal places.)
How does this compare to TargetCo's premerger P/E ratio?
The P/E ratio before the merger was_____ . (Round to two decimal places.)

TargetCo's premerger P/E ratio was__________ . (Round to two decimal places.)
(Select from the drop-down menu.)
Buying TargetCo with stock and creating no synergies, the transaction simply ends-up with a company whose P/E ratio is _______(above/ below/ between) the P/E ratios of the
two companies going into the transaction.
Answer #1

Part (a)

Exchange ratio, E = Price per share of the target / Price per share of the acquirer = 25 /48 = 0.5208

Hence, new shares issued = Shares of target x E = 1.4 mn x 0.5208 = 0.7292 mn

Hence, the EPS after merger = (EPS of Acquirer x Number of shares of acquirer + EPS of target x Number of shares of target) / (Number of shares of acquirer + New shares issued) = (3.96 x 1.2 + 0.99 x 1.4) / (1.2 + 0.7292) = $ 3.18

Part (b)

Exchange ratio, E = Price per share of the target x (1 + premium) / Price per share of the acquirer = 25 x (1 + 20%) /48 = 0.6250

Hence, new shares issued = Shares of target x E = 1.4 mn x 0.6250 = 0.8750 mn

Hence, the EPS after merger = (EPS of Acquirer x Number of shares of acquirer + EPS of target x Number of shares of target) / (Number of shares of acquirer + New shares issued) = (3.96 x 1.2 + 0.99 x 1.4) / (1.2 + 0.8750) = $ 2.96

Part (c)

The correct answer is option C. EPS declines because TargetCo has a higher price-earnings ratio than your firm.

Since you are acquiring a company that has higher P/E ratio than yours, the EPS after merger declines.

--------------------------------

The correct answer is option A. In this case, your shareholders are worse off.

Since EPS post transaction will decline, the shareholders will be worse off.

Part (d)

Price per share Post Merger = ( Market value of acquirer before merger + Market Value of TargetCo. before merger ) / No. of Shares of acquirer post merger

= [(25 x 1.4 Million) + (48 x 1.2 Million)] / 1.9292 million shares = $48

P/E Ratio after the merger = Price per share post merger / EPS Post Merger = $48 / $3.18 = 15.09

-----------------------------------

The P/E ratio before the merger was = Price per share before merger / EPS before Merger = 48 / 3.96 = 12.12

P/E Ratio of TargetCo. before Merger = Price per share before Merger / EPS before Merger = 25 / 0.99 = 25.25

Hence, P/E ratio after the merger is lower than the P/E Ratio of TargetCo. before Merger but higher than the P/E ratio of the acquirer before merger.

------------------------------

Buying TargetCo with stock and creating no synergies, the transaction simply ends-up with a company whose P/E ratio is between the P/E ratios of the two companies going into the transaction.